# On Diverse Preferences for Large Language Model Alignment

### 12 Dec 2023 v1

@@ Abstract
The alignment of large language models (LLMs) with human values is crucial for the development of artificial general intelligence (AGI). One promising approach to achieve this alignment is reinforcement learning from human feedback, which employs a reward model (RM) learned from human preference datasets to guide LLMs in generating text that aligns with human preferences. Through intensive experiments and analysis of reward distribution, this paper finds that preference datasets are diverse from each other, even though they are all proposed to align human preference. Hence, mixing diverse human preference datasets to increase data size for enhancing reward modeling could fail. To address the issue and capture the shared human values from diverse preferences, a new training policy called MORE is introduced, which minimizes preference bias by adaptively adjusting the preference objective across diverse preferences. Experiments with the Pythia-1.4B model and five mixed preference datasets show that MORE achieves superior reward accuracy and lower calibration error, highlighting its ability to leverage diverse human preference data.
@@ 1 Introduction
The emergence of large language models (LLMs) such as ChatGPT and LLama series (Touvron et al., 2023a, b) have propelled us a step forward toward achieving artificial general intelligence (AGI). One key factor in the success of large models is aligning them with human preferences (Christiano et al., 2017), enabling these models to better meet societal expectations, ensuring that their outputs are more useful, benign, and aligned with shared values. Among the alignment algorithms, reinforcement learning from human feedback (RLHF) (Ouyang et al., 2022) has gained prominence as an approach to address the challenge of learning from human preferences. The training process aims to maximize the reward provided by a reward model (RM), which is learned from paired preferred and dispreferred data samples (Bai et al., 2022c; Ganguli et al., 2022; Fernandes et al., 2023). Alternative algorithms to PPO have been continuously proposed (e.g., RLAIF (Lee et al., 2023), ReST (Gulcehre et al., 2023), Rrhf (Yuan et al., 2023), DPO (Rafailov et al., 2023), APO (Cheng et al., 2023b), etc.), bringing to the emergence of numerous outstanding deployable LLMs, such as Anthropic’s Claude (Bai et al., 2022b), Microsoft’s Bing Chat, and Google’s Bard (Manyika, 2023).
The effectiveness of these RLHF series algorithms depends on high-quality RMs (Ouyang et al., 2022). The training of RMs requires a large number of human preference data collected from the real world (typically from crowdworkers) (Bai et al., 2022b). Such human preference data generally needs to be helpful, harmless, benevolent, and unbiased. However, in a pluralistic world, individuals can have different preferences toward the same response due to their diverse experiences, educational backgrounds, religious beliefs, and social values. Moreover, even for the same individual, the value assigned to a specific response from a language model can vary depending on the application scenario (Cheng et al., 2023a). Hence, data annotated by individuals with diverse preferences inherently encompasses a range of opinions and biases. Therefore, leveraging this preference data effectively to train a robust RM poses a significant challenge (Cheng et al., 2023a). The issue arising from these preference divergences or conflicts is of paramount importance, yet it has received little attention and dedicated research thus far in reward modeling.
To identify the challenge, we first investigate a broader range of preferences to understand the inherent characteristics of preference data. We conduct a series of experiments using five open preference datasets, including Helpful&Harmless (Bai et al., 2022a), Webgpt (Nakano et al., 2021a), OASST1 (Köpf et al., 2023), and Summarize (Stiennon et al., 2020). Taking each dataset as a preference with characteristics, we solely train a set of RMs marked by $\{\theta_{\text{helpful}},\theta_{\text{harmless}},\theta_{\text{oasst}},\theta_{\text{webgpt}},\theta_{\text{summ.}}\}$ on each corresponding datasets and testing them on same preference samples as shown in Figure 1. We observe:
  • The implicit preferences across datasets are significantly different. Although these datasets are presented for LLMs to align with human values, the RMs provide very diverse reward values on the same preference data samples. The phenomenon reflects the preference inconsistency hidden in the datasets.
In this paper, we conceptualize the output of RM as a shared reward, accompanied by an additional reward drift. The shared reward indicates the shared human preferences and the reward drift indicates the characteristics. To capture the shared reward in reward modeling, we introduce a straightforward training policy called MORE. The MORE policy aims to train an RM across multiple comparison datasets, enabling the model to capture a broader range of preferences and reduce the impact of reward drift. Our main contributions are encapsulated as follows:
  • We conduct a thorough analysis of how diverse preference datasets affect reward modeling, focusing on the performance of the reward model and the distribution of rewards. Our findings indicate that diverse preference datasets lead to significant variations in reward distributions within the models. Consequently, these variations negatively impact the performance of the reward model.
  • We frame the reward as a shared reward plus reward drift. We propose a straightforward training policy called MORE for reward modeling. This policy only modifies the data sampling and loss definition, adaptively updating the RM by minimizing reward drift.
  • We conduct experiments on the Pythia-1.4B model (Biderman et al., 2023) to evaluate the effects of reward modeling on diverse datasets. Through these experiments, RM trained by MORE achieves the best reward accuracy and the lowest expected calibration error value (Zhu et al., 2023), compared with hybrid and sequential training policies. The results indicate it can output more accurate reward values for the RLHF task.
Figure 2: Overview of MORE policy. The MORE policy consists of the following steps: 1) It collects a diverse batch of data from the mixed dataset. 2) It calculates the gradient about the weights of the reward head. 3) It solves a minimization problem to determine the scalarization factor $\lambda$ for MORE loss. 4) Finally, update the RM via utilized optimizer.
2 Preference Modeling on Diverse Preference Datasets
We study the reward modeling task on diverse preference datasets $\mathcal{D}=\{\mathcal{D}_{1},\dots,\mathcal{D}_{K}\}$. The $K$ denotes the number of diverse preferences and each preference is represented by a set of preference data samples $\mathcal{D}_{i},i\in[K]$. The problem setting covers the case that a single preference dataset implicitly preserves different preferences, such as Helpful&Harmless (Bai et al., 2022a). It also covers the case that mixed different preference datasets were collected in various tasks (Dong et al., 2023), domains (Cheng et al., 2023a), or individuals (Bai et al., 2022b).
2.1 Preliminaries
@@
Preference alignment pipeline. A RM parameterized by $\theta$ can be denoted as a mapping $r_{\theta}:\mathcal{X}\times\mathcal{Y}\rightarrow\mathbb{R}$, which provides a real-valued reward score $r_{\theta}(x,y_{w})$ qualifying a textual response $y=(y_{1},y_{2},\dots,y_{M})\in\mathcal{Y}$ corresponding to an input prompt $x=(x_{1},x_{2},\dots,x_{N})\in\mathcal{X}$. Given a sample $(x,y_{w},y_{l})\sim\mathcal{D}$ from preference dataset $\mathcal{D}$, $r_{\theta}$ is expected to provide a preference score with $r(x,y_{w})>r(x,y_{l})$, indicating the response $y_{w}$ is preferred. The Bradley-Terry model (David, 1963) represents the preference function $p(y_{w}\succ y_{l}|x)$:
$p(y_{w}\succ y_{l}|x)=\frac{\exp{r(x,y_{l})}}{\exp{r(x,y_{l})}+\exp{r(x,y_{w})}},$
where $\sigma(\cdot)$ is an activation usually set as the Sigmoid function. The standard reward modeling process optimizes a pre-trained language model on preference dataset $\mathcal{D}=\{(x_{i},y_{w,i},y_{l,i})\}_{i=1}^{n}$ with the following binary rank loss,
@@
$\mathcal{L}_{\text{rank}}(\theta;\mathcal{D})=-\mathbb{E}_{(x,y_{w},y_{l})\sim\mathcal{D}}\log(\sigma\left(r_{\theta}(x,y_{w})-r_{\theta}(x,y_{l})\right)).$ (1)
Using the reward $r_{\theta}(x,y)$ in the RLHF objective is to optimize for the policy $\pi$ that maximizes the expected reward while minimizing the distance between $\pi$ and some reference policy $\pi_{\text{ref}}$ as:
$\max_{\pi}\mathbb{E}_{x\sim\mathcal{D},y\sim\pi(y|x)}\left[r_{\theta}(x,y)-\beta\mathbb{D}_{\text{KL}}[\pi(y|x)\|\pi_{\text{ref}}(y|x)]\right].$ (2)
2.2 Preference Bias as Reward Drift.
@@
We use $r^{*}(\cdot,\cdot)$ to denote the shared preference reward function, which always gives proper reward value to the contents with the shared values across human beings. As the collected human-feedback datasets are limited and implicitly biased, they usually can not cover the whole shared value. Therefore, training an RM $r_{\theta}$ on a limited preference dataset can be viewed as drifting from an optimal reward. We form the reward drift of a reward model parameterized by $\theta$ from a data level:
$r_{\theta}(x,y)=r_{\theta}^{*}(x,y)+\tilde{r}_{\theta}(x,y),\;\forall(x,y)\in\mathcal{X}\times\mathcal{Y},$ (3)
@@
$\mathcal{L}_{\text{rank}}(\theta;\mathcal{D})$ $=-\mathbb{E}_{(x,y_{w},y_{l})\sim\mathcal{D}}[\log(\sigma(r_{\theta}^{*}(x,y_{w})-r_{\theta}^{*}(x,y_{l})$ (4)
$\quad+\tilde{r}_{\theta}(x,y_{w})-\tilde{r}_{\theta}(x,y_{l})))].$
@@ 2.3 Reward Drift Modeling
We discuss the unique form of learning general preferences from diverse datasets in this part. Firstly, we justify the reward agreement and disagreement in preference modeling:
@@
$r_{\theta}(x,y)=$ $\sum_{i=1}^{K}\lambda_{i}r_{\theta_{i}}(x,y)$ (5)
$=$ $r_{\theta}^{*}(x,y)+\sum_{i=1}^{K}\lambda_{i}\tilde{r}_{\theta_{i}}(x,y),\forall(x,y)\in\mathcal{D},$
where the shared reward $r^{*}_{\theta}(\cdot,\cdot)$ is the same with arbitrary $\lambda$, and $\tilde{r}_{\theta_{i}}(\cdot,\cdot)$ is the reward drift. We interpret that the $\tilde{r}_{\theta_{i}}(\cdot,\cdot)$ is provided by subset of parameters $\theta_{i}$, representing the preferences from the $i$-th dataset $\mathcal{D}_{i}$. For example, the $\theta$ can be implemented as an ensemble model, where $\{{\theta_{i}}\},i\in[N]$ is the base models. Therefore, it is natural to adjust the weight $\lambda$ in an ensemble manner (Coste et al., 2023; Jang et al., 2023; Touvron et al., 2023a) to mitigate the reward drift such that $\min\sum_{i=1}^{K}\lambda_{i}\tilde{r}_{\theta_{i}}(x,y)$. Moreover, if the datasets are extremely diverse, the data samples are very likely to be disagreed by $K$ diverse preference by Definition 1. Hence, there always exists an combination $\lambda$ such that $\sum_{i=1}^{K}\lambda_{i}\tilde{r}_{\theta_{i}}(x,y)=0$.
2.4 Training Policy: MORE
Compared with an ensemble of a set of RMs (Jang et al., 2023), we are more interested in training a single RM that learns the shared preference. Our idea is to reduce the model update on reward drift during RM training via linear scalarization (Barrett & Narayanan, 2008).
@@ 2.4 Training Policy: MORE
$\mathcal{L}_{\text{MORE}}(\theta;\mathcal{D})=\sum_{i=1}^{K}\lambda_{i}\mathcal{L}_{\text{rank}}(\theta,\mathcal{D}_{i}),$ (6)
@@
Computing $\lambda$. We propose to use adaptive weighting methods to balance the bias across different domains and adjust the reward modeling process. Straightforwardly, the mitigation task in Equation 9 can be efficiently solved by the Frank-Wolfe solver (Jaggi, 2013; Sener & Koltun, 2018; Zhou et al., 2022; Zeng et al., 2023), while the computing cost for solving it is proportional to the size of parameters $\theta$. To avoid expensive computation, we only utilize gradients of the reward head $\theta_{\text{rm}}\in\mathcal{R}^{h}$ from each preference. In detail, we obtain the hidden states $z=r_{\theta_{\text{lm}}}(x)$ before the reward head and compute the gradient of the reward head solely with data $(z,y_{w},y_{l})$. In summary, the $\lambda$ is computed by:
$\lambda=\arg\min_{\lambda}\left\|\sum_{i=1}^{K}\lambda_{i}g_{i}\right\|^{2},s.t.,\;\sum_{i=1}^{K}\lambda_{i}=1,\lambda_{i}\geq 0,$ (7)
@@
Outline. The MORE loss requires simple modification on batch data sampling and additional computing $\lambda$. In detail, the pipeline of MORE is given in Figure 2 and described as follows: 1) Sample a diverse batch data =$\mathcal{B}=\{\mathcal{B}_{i}\}_{i=1}^{K}$, $\mathcal{B}_{i}=\{x,y_{w},y_{l}\}_{b=1}^{B}$ and input the batch data forward the RM and obtain the hidden states $\{z_{i}\}_{b=1}^{B}$ before the reward head $\theta_{\text{rm}}$. 2) Compute the gradient of reward with data $\{z,y_{w},y_{l}\}$ and update the cache $\{g_{i}^{t}\}$ following equation 8. 3) Compute the weights $\lambda$ by equation 7. Finally, we substitute the loss weights in equation 6 as the final loss for the optimizer to conduct backward and model updating. This procedure mitigates the potential reward drift from the overall reward.
@@
3 Experiments
Table 1: Reward model performance on diverse datasets.

### 25 Dec 2023 v2

# On Diversified Preferences of Large Language Model Alignment

@@ Abstract
Aligning large language models (LLMs) with human preferences has been recognized as the key to improving LLMs’ interaction quality. However, in this pluralistic world, human preferences can be diversified by people’s different tastes, which hinders the effectiveness of LLM alignment methods. In this paper, we provide the first quantitative analysis to verify the existence of diversified preferences in commonly used human feedback datasets. To mitigate the alignment ineffectiveness caused by diversified preferences, we propose a novel Multi-Objective Reward learning method (MORE), which can automatically adjust the learning gradients across different preference data sources. In experiments, we evaluate MORE with the Pythia-1.4B model on five mixed human preference datasets, on which our method achieves superior performance compared with other baselines in terms of preference accuracy and prediction calibration.
@@ 1 Introduction
Large language models (LLMs), such as ChatGPT (OpenAI, 2023) and LLaMA (Touvron et al., 2023a, b), have significantly accelerated the development process toward artificial general intelligence (AGI). Among the key factors for such great achievement, the alignment technique, which finetunes LLMs with human feedback (Christiano et al., 2017), has played an essential role in training LLMs’ responses to follow human values (e.g., helpfulness and harmlessness) (Askell et al., 2021). Among the LLM alignment algorithms, reinforcement learning from human feedback (RLHF) (Ouyang et al., 2022) has become the mainstream solution, which first learns a reward model (RM) representing human preferences and then updates LLMs via the proximal policy optimization (PPO) (Schulman et al., 2017) toward generating responses with higher RM scores. However, RLHF has recently been challenged for its cumbersome computational complexity and training instability. To avoid the reinforcement learning scheme of RLHF, alternative alignment methods have been sequentially proposed, such as RAFT (Dong et al., 2023b), DPO (Rafailov et al., 2023), RRHF (Yuan et al., 2023), and APO (Cheng et al., 2023b).
@@
The performance of these LLM alignment methods highly depends on the quality of human preference data $(x,y_{w},y_{l})$, where $x$ is the input query to the LLM, and response $y_{w}$ is preferred to response $y_{l}$ under the human annotation (Ouyang et al., 2022). Ideally, the preference data sets should uniformly be helpful, harmless, benevolent, and unbiased to guide the LLM alignment. However, in real-world scenarios, individuals can have diversified preferences on the same topic based on their different experiences, educational backgrounds, religions, and cultures. Even for the same person, his or her expected model answer to a particular question can vary depending on different scenarios (Cheng et al., 2023a). This causes a serious inconvenience in preference data annotation and collection processes, for different annotators can make considerable disagreements on the selection of better responses (Bai et al., 2022a). The annotation disagreement significantly hinders the effectiveness of alignment methods (Davani et al., 2022; Wan et al., 2023). Hence, how to resolve divergences and conflicts in preference data is vitally important, yet it has not received adequate attention and remained unexplored in the LLM alignment domain.
[…]
To address the problem of diversified preferences, inspired by multi-objective optimization methods (Sener & Koltun, 2018; Zeng et al., 2023), we regard RMs as a shared reward additionally with a customized reward drift. The shared reward represents the shared human preferences (or general preferences) and the reward drift contains individual or domain-specific preference information (Cheng et al., 2023a). To capture the shared (general) preference information, we introduce a Multi-Objective Reward training scheme (MORE), which treats feedback data from different sources as different learning objectives, then re-weight these objectives by minimizing the mean gradient of their corresponding reward drifts. With multi-objective reward learning, RMs can capture a broader range of preferences and mitigate the impact of preference conflicts. The main contributions of this paper are:
  • We conduct a comprehensive analysis of how diversified preferences affect reward modeling in LLM alignment. We find that diversified preferences can cause high variances in the RM output distributions, which negatively impact the performance of LLM alignment.
  • We model RMs as a shared reward plus reward drift and propose a Multi-Objective Reward (MORE) training scheme, which makes self-adaption to the RM learning gradient by minimizing the reward drifts.
  • We empirically verified the effectiveness of MORE with Pythia-1.4B (Biderman et al., 2023) RMs on multiple preference datasets, where MORE outperforms other baseline training strategies for multi-source preference learning and alignment.
Figure 2: Multi-objective reward model training scheme (MORE), which consists of four steps: (1) collect a diversified batch of data from the mixed dataset; (2) calculate the RM gradient for each preference source; (3) minimize the reward drift to determine the scalar $(\lambda_{1},\lambda_{2},\dots,\lambda_{K})$ for MORE loss; (4) update the RM with the re-weighted RM loss.
2 Preliminary
Large Language Model Alignment
@@ 2 Preliminary
Parameterized by $\theta$, a reward model (RM) is a mapping $r_{\theta}:\mathcal{X}\times\mathcal{Y}\rightarrow\mathbb{R}$, which provides a real-valued reward score $r_{\theta}(x,y)$ evaluating a textual response $y=(y_{1},y_{2},\dots,y_{M})\in\mathcal{Y}$ corresponding to an input prompt $x=(x_{1},x_{2},\dots,x_{N})\in\mathcal{X}$. Given a sample $(x,y_{w},y_{l})\sim\mathcal{D}$ from a preference dataset $\mathcal{D}$, $r_{\theta}$ is expected to provide a preference score with $r_{\theta}(x,y_{w})>r_{\theta}(x,y_{l})$, representing the response $y_{w}$ is preferred. Following the Bradley-Terry model (David, 1963), the RM learning objective on on the preference dataset $\mathcal{D}=\{(x,y_{w},y_{l})\}$ is defined as:
@@
$\mathcal{L}_{\text{rank}}(\theta;\mathcal{D})=-\mathbb{E}_{\mathcal{D}}\log(\sigma\left(r_{\theta}(x,y_{w})-r_{\theta}(x,y_{l})\right))$ (1)
where $\sigma(\cdot)$ is the Sigmoid function. With a well-learned reward $r_{\theta}(x,y)$, LLM alignment optimizes the generation policy $\pi_{\phi}(y|x)$ by maximizing the expected reward value:
$\mathbb{E}_{x\sim\mathcal{D},y\sim\pi(y|x)}[r_{\theta}(x,y)]-\beta\mathbb{D}_{\text{KL}}[\pi(y|x)\|\pi_{\text{ref}}(y|x)],$ (2)
where $\mathbb{D}_{\text{KL}}[\pi(y|x)\|\pi_{\text{ref}}(y|x)]$ is the KL divergence regularizer between current policy $\pi_{\phi}$ and a reference $\pi_{\text{ref}}$, preventing the optimization from instability and degeneration.
Domain Generalization
Machine learning methods suffer from performance degeneration when the source domain data and the target domain data follow different distributions, which has been recognized as the domain shift problem (Pan & Yang, 2009; Csurka, 2017; Wang et al., 2021). To address this problem, domain generalization is proposed to minimize the domain shift across domains. In this direction, existing methods aim to learn the domain invariant representation to reduce the discrepancy between representations of multiple source domains (Zhou et al., 2022a).
Multi-objective Optimization
Multi-objective Optimization (MOO) (Gunantara, 2018) is a branch of methods addressing learning problems involving multiple conflicting objectives. In real-world scenarios, it commonly encounters situations where multiple objectives need to be considered simultaneously, often with trade-offs between them. In the practice of machine learning, most MOO methods (Sener & Koltun, 2018; Zeng et al., 2023) apply linear scalarization (Barrett & Narayanan, 2008) to merge multiple objectives into one, and then automatically adjust the objective coefficients to balance the conflicts among different tasks.
3 Empirical Study of Diversified Preferences
We start with an empirical analysis of diversified preferences in reward modeling on multiple sources $\mathcal{D}=\{\mathcal{D}_{1},\dots,\mathcal{D}_{K}\}$, where each data source $D_{k}$ contains the preference comparison pairs from different tasks (Dong et al., 2023a), domains (Cheng et al., 2023a), or individuals (Bai et al., 2022a). As mentioned in Section 1, we selected Summarize (Stiennon et al., 2020), Webgpt (Nakano et al., 2021a), Helpful&Harmless (Bai et al., 2022a), and OASST1 (Köpf et al., 2023) as the different preference sources to empirical analysis the phenomena of diversified preferences. We use Pythia-1.4B (Biderman et al., 2023) as the RM base, and fine-tuned RMs with comparisons from each source. The learning rate is set to $5e^{-5}$. All RMs are trained with batch size 64 in one epoch.
The reward distribution of different reward models is diversified on the same dataset. We summarize the reward difference (the reward difference is the difference in reward value between the winning sample and the losing sample given by a reward model) provided by RMs, as shown in Figure 3. Additional results on other test sets are provided in Appendix C. Compared with the results of raw model $\text{RM}_{\text{Raw}}$, we observe that training on different datasets results in diverse reward values (right plot) and distribution shift (left plot). Specifically, the reward value distribution of $\text{RM}_{\text{Harmless}}$ shifts from the $\text{RM}_{\text{Raw}}$ in a certain degree. While the reward value distributions of $\text{RM}_{\text{Helpful}}$, $\text{RM}_{\text{Webgpt}}$, $\text{RM}_{\text{Oasst1}}$ and $\text{RM}_{\text{Summ.}}$ shifts to the a different direction. Moreover, despite of the distribution of $\text{RM}_{\text{Helpful}}$, $\text{RM}_{\text{Webgpt}}$, $\text{RM}_{\text{Oasst1}}$ and $\text{RM}_{\text{Summ.}}$ are similar, the mean-variance of their reward values are quite different.
Furthermore, when considering the accuracy gains illustrated in Figure 1, the observed shift in reward distribution indicates that the learned reward values from preference datasets are diversified. To effectively capture the shared reward values across these diversified preferences, it becomes necessary to formulate a new problem approach for reward modeling on diverse preference datasets.
Figure 3: Statistics of reward values on Harmless test set. Left: the distribution of reward differences. The box plot depicts the statistics of reward differences on the test set. Right: The solid box indicates the statistics of reward values on correct rewarded samples, while the hollow box indicates the wrong rewarded samples.
4 Multi-objective Reward Learning
4.1 Preference Diversity as Reward Drift
@@
We denote $r^{*}(\cdot,\cdot)$ as the shared reward function, which (ideally) provides reward values reflecting the shared values among people. As the collected human-feedback datasets are limited and implicitly biased, they usually can not cover the whole shared value. Therefore, training an RM $r_{\theta}$ on a limited preference dataset can be viewed as drifting from an optimal reward. We form the reward drift of a reward model parameterized by $\theta$ from a data level:
$r_{\theta}(x,y)=r_{\theta}^{*}(x,y)+\tilde{r}_{\theta}(x,y)$ (3)
@@
$\mathcal{L}_{\text{rank}}(\theta;\mathcal{D})=-$ $\mathbb{E}_{(x,y_{w},y_{l})\sim\mathcal{D}}[\log(\sigma(r_{\theta}^{*}(x,y_{w})-r_{\theta}^{*}(x,y_{l})$
$\quad+\tilde{r}_{\theta}(x,y_{w})-\tilde{r}_{\theta}(x,y_{l})))].$ (4)
@@ 4.2 Reward Drift Modeling
We discuss the unique form of learning general preferences from diversified data sources in this section. Firstly, we justify the reward agreement and disagreement in preference modeling:
@@
$r_{\theta}(x,y)=\sum_{i=1}^{K}\lambda_{i}r_{\theta_{i}}(x,y)=r_{\theta}^{*}(x,y)+\sum_{i=1}^{K}\lambda_{i}\tilde{r}_{\theta_{i}}(x,y)$ (5)
where the shared reward $r^{*}_{\theta}(\cdot,\cdot)$ is the same with arbitrary $\lambda$, and $\tilde{r}_{\theta_{i}}(\cdot,\cdot)$ is the reward drift. We interpret that the $\tilde{r}_{\theta_{i}}(\cdot,\cdot)$ is provided by subset of parameters $\theta_{i}$, representing the preferences from the $i$-th dataset $\mathcal{D}_{i}$. For example, the $\theta$ can be implemented as an ensemble model, where $\{{\theta_{i}}\},i\in[N]$ is the base models. Therefore, it is natural to adjust the weight $\lambda$ in an ensemble manner (Coste et al., 2023; Jang et al., 2023; Touvron et al., 2023a; Eisenstein et al., 2023) to mitigate the reward drift such that $\min\sum_{i=1}^{K}\lambda_{i}\tilde{r}_{\theta_{i}}(x,y)$. Moreover, if the datasets are extremely diverse, the data samples are very likely to be disagreed by $K$ diverse preference by Definition 1. Hence, there always exists an combination $\lambda$ such that $\sum_{i=1}^{K}\lambda_{i}\tilde{r}_{\theta_{i}}(x,y)=0$.
4.3 Training Scheme: MORE
Compared with an ensemble of a set of RMs (Jang et al., 2023; Eisenstein et al., 2023), we are more interested in training a single RM that learns the shared preference. Our idea is to reduce the model update on reward drift during RM training via linear scalarization (Barrett & Narayanan, 2008).
@@ 4.3 Training Scheme: MORE
$\textstyle\mathcal{L}_{\text{MORE}}(\theta;\mathcal{D})=\sum_{i=1}^{K}\lambda_{i}\mathcal{L}_{\text{rank}}(\theta,\mathcal{D}_{i}),$ (6)
@@
Computing $\lambda$
We propose to use adaptive weighting methods to balance the bias across different domains and adjust the reward modeling process. Straightforwardly, the mitigation task in Equation 9 can be efficiently solved by the Frank-Wolfe solver (Jaggi, 2013; Sener & Koltun, 2018; Zhou et al., 2022b; Zeng et al., 2023), while the computing cost for solving it is proportional to the size of parameters $\theta$. To avoid expensive computation, we only utilize gradients of the reward head $\theta_{\text{rm}}\in\mathcal{R}^{h}$ from each preference. In detail, we obtain the hidden states $z=r_{\theta_{\text{lm}}}(x)$ before the reward head and compute the gradient of the reward head solely with data $(z,y_{w},y_{l})$. In summary, the $\lambda$ is computed by:
$\lambda=\arg\min_{\lambda}\left\|\sum_{i=1}^{K}\lambda_{i}g_{i}\right\|^{2},\ \sum_{i=1}^{K}\lambda_{i}=1,\lambda_{i}\geq 0,$ (7)
@@
Outline
The MORE loss requires simple modification on batch data sampling and additional computing $\lambda$. In detail, the pipeline of MORE is given in Figure 2 and described as follows: 1) Sample a diverse batch data =$\mathcal{B}=\{\mathcal{B}_{i}\}_{i=1}^{K}$, $\mathcal{B}_{i}=\{x,y_{w},y_{l}\}_{b=1}^{B}$ and input the batch data forward the RM and obtain the hidden states $\{z_{i}\}_{b=1}^{B}$ before the reward head $\theta_{\text{rm}}$. 2) Compute the gradient of reward with data $\{z,y_{w},y_{l}\}$ and update the cache $\{g_{i}^{t}\}$ following equation 8. 3) Compute the weights $\lambda$ by equation 7. Finally, we substitute the loss weights in equation 6 as the final loss for the optimizer to conduct backward and model updating. This procedure mitigates the potential reward drift from the overall reward.
@@
Table 1: Reward model performance on diverse datasets. $*{}^{*}$ower is better.